A 1.79-g sample of an unknown gas has a volume of 530 mL and a pressure of 923 mm Hg at 36.6 °C. Calculate the molar mass of this compound.

________ g/mol
1
Expert's answer
2019-03-19T06:02:29-0400

A 1.79 g sample of an unknown gas has a volume of 530 mL and a pressure of 923 mm Hg at 36.6 °C. Calculate the molar mass of this compound.

m(gas) = 1.79 g

V(gas) = 530 ml = 0.00053 m3

p(gas) = 923 mm Hg  = 123056.5 Pa

T = 36.6°C = 309.6 K

R = 8.31 J⋅mol−1⋅K−1

M(gas) -?

Solution:

"p*V = (m*R*T)\/M""M = (m*R*T)\/ (p*V)"

M(gas) = (1.79 g*8.31 J⋅mol−1⋅K−1*309.6 K) / (123056.5 Pa*0.00053 m3) = 70.61 g/mol

Answer:

The molar mass of this compound is 70.61 g/mol.